Purple is 1/4 of the wheel, Cyan is 2/5 and the rest of the wheel is Fuscia. If you land on purple you win $40, Cyan you win $11 and otherwise you win nothing. The game costs $17 to play. What is your expected value per play (after you pay the cost to play)?
Found 2 solutions by Boreal, greenestamps:
Answer by Boreal(15235)

E(x)=sum x*p(x)
=40(1/4)+11(2/5)-17(1); since one always has to pay $17 to play.
=10+4.4-17
=-$2.60

Answer by greenestamps(13216)



P(purple)=0.25
P(cyan)=0.40
P(fuschia)=1-(0.25+0.40)=1-0.65=0.35

expected value for purple: 0.25(-17+40)=0.25(23)=+5.75
expected value for cyan: 0.40(-17+11)=0.40(-6)=-2.40
expected value for fuschia: 0.35(-17+0)=0.35(-17)=-5.95

expected value per play: 5.75-2.4-5.95 = -2.6